OpenStudy (anonymous):

Use gauss approach to find the following sum (do not use formulas) A. 1+3+5+7+...+101

OpenStudy (anonymous):

there is a simple formula for both of these, but i think "gauss approach" means add forward and backward, then divide by 2 \[S=1 + 2+ 3+...+1000+ 1001+1002\] \[S=1002+1001+1000+...+ 3 + 2 +1\] \[2S=1003+1003+1003+...+1003\] \[2S=1002\times 1003\] \[S=\frac{1002\times 1003}{2}\]

OpenStudy (campbell_st):

gauss's approach is to reverse the order then add the vertical pair e.g 1 2 3 4 5 6 + 6 5 4 3 2 1 --------------------- 7 7 7 7 7 7 the vertical pairs will always add to the same value to you have 6 lots of 7 so 6 x 7 = 42 and becuase you have 2 lots of the numbers 1 to 6 you need to halve the answer 42/2 = 21 so the sum of the 1st 6 numbers is 21

OpenStudy (anonymous):

from this derive the formula \[\sum_{k=1}^nk=\frac{n(n+1)}{2}\]

OpenStudy (campbell_st):

so in your question for the sum of the numbers 1 2 3 4 ........ 1002 +1002 1001 1000 999 1 ------------------------------------ 1003 1003 1003 1003 1003 you have 1002 lots of 1003 then halve that answer